Harvey noticed that his gas is 1/4 full. He completely filled the tank by putting 9 gallons into it. What is the capacity of the gas tank?
Answer by RAY100(1637)

Let x be capacity of full tank
.
3/4*x = 9,,,,,,,multiply by 4/3
.
x= 12 gallons
.
check ,,,,,,9/12=3/4,,,,ok
.